OpenStudy (anonymous):

What conic section does the equation 9x^2 + 4y^2 + 36x = 0 represent? circle ellipse hyperbola parabola

OpenStudy (anonymous):

ellipse..

OpenStudy (anonymous):

9(x+2)^2+4y^2=36 (x+2)^2/4 +y^2/9=1

OpenStudy (anonymous):

The equation of a conic section is 2y^2 - x - 8y + 5 = 0. What is the equation in graphing form? Option A: x - 1 = 2(y - 4)^2 Option B: x + 3 = 2(y - 2)^2 Option C: x^2/2 +(y-4)^2/4 = 1 Option D: x^2/2 - (y+5^2)/4 = 1